### v3

@@ Abstract
This paper introduces deep super ReLU networks (DSRNs) as a method for approximating functions in Sobolev spaces measured by Sobolev norms $W^{m,p}$ for $m\in\mathbb{N}$ with $m\geq 2$ and $1\leq p\leq+\infty$. Standard ReLU deep neural networks (ReLU DNNs) cannot achieve this goal. DSRNs consist primarily of ReLU DNNs, and several layers of the square of ReLU added at the end to smooth the networks output. This approach retains the advantages of ReLU DNNs, leading to the straightforward training. The paper also proves the optimality of DSRNs by estimating the VC-dimension of higher-order derivatives of DNNs, and obtains the generalization error in Sobolev spaces via an estimate of the pseudo-dimension of higher-order derivatives of DNNs.
@@ 1 Introduction
Deep neural networks (DNNs) with the rectified linear unit (ReLU) activation function [13] have become increasingly popular in scientific and engineering applications, including image classification [20, 18], regularization [6], and dynamic programming [12, 39]. However, when tasked with approximating higher-order derivatives of target functions, such as in solving partial differential equations (PDEs) [21, 8, 31, 10] by DNNs, ReLU DNNs are unable to provide accurate results due to their lack of smoothness. For instance, physics-informed neural networks (PINNs) [31] have emerged as a popular approach to solving PDEs. When solving a PDE defined on a domain $\Omega=[a,b]^{d}$
$\begin{cases}\Delta u=f&\text{ in }\Omega,\\ u=g&\text{ on }\partial\Omega,\end{cases}$
the loss functions $L(\boldsymbol{\theta})$ is defined as
$L(\boldsymbol{\theta}):=\|\Delta\phi(\boldsymbol{x};\boldsymbol{\theta})-f\|^{2}_{L^{2}(\Omega)}+\beta\|\phi(\boldsymbol{x};\boldsymbol{\theta})-g\|^{2}_{L^{2}(\partial\Omega)},$
where $\phi(\boldsymbol{x};\boldsymbol{\theta})$ denotes a deep neural network (DNN) with parameters $\boldsymbol{\theta}$, and $\beta>0$ is a constant that balances the importance of the boundary and interior. However, ReLU DNNs are unsuitable for solving such PDEs because they lack the smoothness required to accurately calculate the Laplacian of $\phi(\boldsymbol{x};\boldsymbol{\theta})$, i.e., $\Delta\phi(\boldsymbol{x};\boldsymbol{\theta})$ does not exist.
@@
The approximation by shallow neural networks has been explored in [28, 29, 25, 40, 9, 35] and [27]. Regarding the approximation of deep neural networks (DNNs) measured by $L^{p}$-norms where $1\leq p\leq\infty$, various studies have investigated this topic, as seen in [26, 32, 33, 22, 5] and [23].The Sobolev norm approximation of DNNs has been the subject of study in several works, including [14]; [15]; [19]; [41] and [42]. In [14], ReLU DNNs were used to approximate functions measured by $W^{s,p}$ where $s\leq 1$, but their method cannot be generalized to $s>1$ due to the lack of smoothness of ReLU. Smooth activation functions were applied in [15], which achieved a non-optimal approximation rate, and training with these functions may be difficult due to the complexity of activation functions. In [19] and [42], ReLU and its squared version were utilized as DNN activation functions, yet the following three questions remain.
@@
Firstly, the problem of approximating Sobolev functions with DNNs exhibiting super convergence rates, measured by $W^{m,p}$ norms for $p\in[1,\infty)$, remains unresolved, although the $W^{m,\infty}$ case was addressed by [42], and the case of $m=0$ was considered by [34]. Secondly, proper distributions of the squared ReLU function remain unknown. Incorporating it at any layer of a DNN with a depth of $L$ might lead to a polynomial degree of $2^{L}$ with respect to parameters, which poses challenges in training. Thirdly, the optimality of the approximation process remains unclear due to the complexity associated with higher-order derivatives of DNNs using ReLU and squared ReLU activations. To establish optimality, estimating the optimal Vapnik-Chervonenkis dimension (VC-dimension) and pseudo-dimension [2, 37, 1, 30] of their higher order derivatives, which characterize the complexity or richness of a function set, need to be estimated. Even for DNNs without derivatives, as demonstrated in [3], obtaining optimal bounds is challenging. The upper VC-dimension and pseudo-dimension for DNNs using squared ReLU is $\mathcal{O}(N^{2}L^{3}\log_{2}N\log_{2}L)$, where $L$ signifies DNN depth and $N$ denotes DNN width. However, this result is not directly applicable to proving the optimality of squared ReLU-based DNN approximation due to a polynomial gap in depth $L$.
@@
In this paper, the deep super ReLU networks (DSRNs) is presented to address these three questions, as illustrated in Figure 1. The DSRN architecture consists of two components for a network with depth $L$: ReLU DNNs with a depth of approximately $\mathcal{O}(L)$, and the square of ReLU DNNs with a depth of approximately $\mathcal{O}(\log_{2}L)$—significantly shallower than the first component.
@@
In terms of training, our DSRNs predominantly employ ReLU activations, with the square of ReLU appearing only at the end of the DNNs. This enhances the smoothness of DNNs while preserving a straightforward network structure without intricate composition formulas. This contrasts with [42], which did not specify where to apply the square of ReLU, leading to training difficulties. Our experiments in Sec. 7 demonstrate that our proposed DNN (DSRN) trains effectively, unlike those using the square of ReLU throughout, which are difficult to train due to gradient vanishing.
Finally, we estimate the generalization error of DSRNs using two methods. One method (Theorem 6.1) provides a better rate with respect to the number of parameters, while the other method (Theorem 6.2) achieves a better rate concerning the number of sample points but has a relative worse rate regarding the number of parameters. Both methods apply estimates of the VC-dimension and pseudo-dimension of higher-order derivatives of deep neural networks. To our knowledge, this is the first work to consider the VC-dimension and pseudo-dimension of higher-order derivatives in neural networks, which can be directly applied in the error estimation of Sobolev training.
@@
$\bullet$ We provide an upper bound estimation for the pseudo-dimension of high-order derivatives of DNNs. By utilizing this estimation, we can obtain the generalization error of DNNs in Sobolev spaces that contain high-order derivatives of DNNs.
@@ 2.1 Neural networks
1. Matrices are denoted by bold uppercase letters. For example, $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ is a real matrix of size $m\times n$ and $\boldsymbol{A}^{\mathrm{T}}$ denotes the transpose of $\boldsymbol{A}$.
2. Vectors are denoted by bold lowercase letters. For example, $\boldsymbol{v}\in\mathbb{R}^{n}$ is a column vector of size $n$. Furthermore, denote $\boldsymbol{v}(i)$ as the $i$-th elements of $\boldsymbol{v}$.
3. For a $d$-dimensional multi-index $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\cdots\alpha_{d}]\in\mathbb{N}^{d}$, we denote several related notations as follows:
$(a)~{}|\boldsymbol{\alpha}|=\left|\alpha_{1}\right|+\left|\alpha_{2}\right|+\cdots+\left|\alpha_{d}\right|;$
$(b)~{}\boldsymbol{x}^{\boldsymbol{\alpha}}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{d}^{\alpha_{d}},~{}\boldsymbol{x}=\left[x_{1},x_{2},\cdots,x_{d}\right]^{\mathrm{T}};$
$(c)~{}\boldsymbol{\alpha}!=\alpha_{1}!\alpha_{2}!\cdots\alpha_{d}!.$
4. Assume $\boldsymbol{n}\in\mathbb{N}_{+}^{m}$, and $f$ and $g$ are functions defined on $\mathbb{N}_{+}^{m}$, then $f(\boldsymbol{n})=\mathcal{O}(g(\boldsymbol{n}))$ means that there exists positive $C$ independent of $\boldsymbol{n},f,g$ such that $f(\boldsymbol{n})\leq Cg(\boldsymbol{n})$ when all entries of $\boldsymbol{n}$ go to +$\infty$+ ∞.
5. Define $\sigma_{1}(x):=\sigma(x)=\max\{0,x\}$ and $\sigma_{2}:=\sigma^{2}(x)$. We call the neural networks with activation function $\sigma_{t}$ with $t=i$ as $\sigma_{i}$ neural networks ($\sigma_{i}$-NNs), $i=1,2$. With the abuse of notations, we define $\sigma_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ as $\sigma_{i}(\boldsymbol{x})=\left[\begin{array}[]{c}\sigma_{i}(x_{1})\\ \vdots\\ \sigma_{i}(x_{d})\end{array}\right]$ for any $\boldsymbol{x}=\left[x_{1},\cdots,x_{d}\right]^{T}\in\mathbb{R}^{d}$.
@@
$\boldsymbol{x}=\tilde{\boldsymbol{h}}_{0}\stackrel{{\scriptstyle W_{1},b_{1}}}{{\longrightarrow}}\boldsymbol{h}_{1}\stackrel{{\scriptstyle\sigma_{i}}}{{\longrightarrow}}\tilde{\boldsymbol{h}}_{1}\ldots\stackrel{{\scriptstyle W_{L},b_{L}}}{{\longrightarrow}}\boldsymbol{h}_{L}\stackrel{{\scriptstyle\sigma_{i}}}{{\longrightarrow}}\tilde{\boldsymbol{h}}_{L}\stackrel{{\scriptstyle W_{L+1},b_{L+1}}}{{\longrightarrow}}\phi(\boldsymbol{x})=\boldsymbol{h}_{L+1},$
where $\boldsymbol{W}_{i}\in\mathbb{R}^{N_{i}\times N_{i-1}}$ and $\boldsymbol{b}_{i}\in\mathbb{R}^{N_{i}}$ are the weight matrix and the bias vector in the $i$-th linear transform in $\phi$, respectively, i.e., $\boldsymbol{h}_{i}:=\boldsymbol{W}_{i}\tilde{\boldsymbol{h}}_{i-1}+\boldsymbol{b}_{i},~{}\text{ for }i=1,\ldots,L+1$ and $\tilde{\boldsymbol{h}}_{i}=\sigma_{i}\left(\boldsymbol{h}_{i}\right),\text{ for }i=1,\ldots,L.$ In this paper, an DNN with the width $N$ and depth $L$, means (a) The maximum width of this DNN for all hidden layers is less than or equal to $N$. (b) The number of hidden layers of this DNN is less than or equal to $L$.

### v4

@@ Abstract
This paper introduces deep super ReLU networks (DSRNs) as a method for approximating functions in Sobolev spaces measured by Sobolev norms $W^{m,p}$ for $m\in\mathbb{N}$ with $m\geq 2$ and $1\leq p\leq+\infty$. Standard ReLU deep neural networks (ReLU DNNs) cannot achieve this goal. DSRNs consist primarily of ReLU DNNs, and several layers of the square of ReLU added at the end to smooth the networks output. This approach retains the advantages of ReLU DNNs, leading to the straightforward training. The paper also proves the optimality of DSRNs by estimating the Vapnik–Chervonenkis dimension of higher-order derivatives of DNNs, and obtains the generalization error in Sobolev spaces via an estimate of the pseudo dimension of higher-order derivatives of DNNs.
keywords:
ReLU nueral network; Squared ReLU nueral network; Sobolev spaces; Vapnik–Chervonenkis dimension; pseudo dimension.
@@ 1 Introduction
Deep neural networks (DNNs) with the rectified linear unit (ReLU) activation function [20] have become increasingly popular in scientific and engineering applications, including image classification [33, 27], regularization [8], and dynamic programming [19, 57]. However, when tasked with approximating higher-order derivatives of target functions, such as in solving partial differential equations (PDEs) [34, 14, 48, 16] by DNNs, ReLU DNNs are unable to provide accurate results due to their lack of smoothness. For instance, physics-informed neural networks (PINNs) [48] have emerged as a popular approach to solving PDEs. When solving a PDE on the domain $\Omega=[a,b]^{d}$,
$\begin{cases}\mathcal{L}u=f&\text{in }\Omega,\\ u=g&\text{on }\partial\Omega,\end{cases}$
the loss function $L(\bm{\theta})$ is defined as
$L(\bm{\theta}):=\|\mathcal{L}\phi(\bm{x};\bm{\theta})-f\|_{L^{2}(\Omega)}^{2}+\beta\,\|\phi(\bm{x};\bm{\theta})-g\|_{L^{2}(\partial\Omega)}^{2},$
where $\phi(\bm{x};\bm{\theta})$ denotes a deep neural network (DNN) with parameters $\bm{\theta}$, $\beta>0$ is a weight balancing the boundary and interior terms, and $\mathcal{L}$ is a second-order differential operator. However, ReLU-based DNNs are unsuitable for this task because they lack the smoothness required to compute $\mathcal{L}\phi(\bm{x};\bm{\theta})$ accurately; in fact, $\mathcal{L}\phi(\bm{x};\bm{\theta})$ may not even exist in the classical sense.
@@
The approximation by neural networks in various spaces measured by variable norms has been established in a lot of papers. The following table (Table 1) lists the references and shows the difference between their work and ours.
Measured Norm / Target Space Activations Architecture References
$L^{p}$/Sobolev spaces G.A [^1] G.A (general activations) indicates a broad activation class: either multiple specific activations are treated, or a generalized activation satisfying stated conditions. Shallow [41, 36, 60]
$^{2}, LL^{\infty}$, $H^{1}$
/(spectral) Barron spaces
G.A [15, 53, 54, 32]
$L^{p}$/Smooth spaces ReLU Deep [50, 51, 38, 46]
$L^{p}$/Sobolev spaces ReLU and $\tanh$ [61, 52, 11]
$L^{p}$ and $H^{1}$
/Special spaces
ReLU or RePU
[43, 55, 5, 58]
[44, 45]
$L^{p}$/Smooth spaces G.A [62]
$W^{\alpha,p}$, $0\leq\alpha\leq 1$
/Sobolev spaces
ReLU [59, 21]
$W^{m,\infty}$$W^{n,\infty}$ G.A rate $\epsilon^{-\frac{d}{n-m}}$ [22, 26, 25]
$W^{m,p}$$W^{n,p}$
ReLU and
squared ReLU
rate $\epsilon^{-\frac{d}{2(n-m)}}$ Theorem 2 and Corollary 8
Table 1: Summary of approximation results across various norms/spaces, activation choices, and architectures for fully connected neural networks. In the last two rows, the rate indicates the number of parameters required for a neural network to achieve an $\epsilon$ approximation accuracy measured in the specified norms.
The above table summarizes only a portion of the approximation results for fully connected neural networks. Other network architectures, such as convolutional neural networks [63, 18, 40] and transformers [7, 24], are not discussed here, as they are not closely related to the focus of our work.
In [59], ReLU and its squared version were utilized as DNN activation functions, yet the following three questions remain.
@@
Firstly, the problem of approximating Sobolev functions with DNNs exhibiting super convergence rates, measured by $W^{m,p}$ norms for $p\in[1,\infty)$, remains unresolved, although the $W^{m,\infty}$ case was addressed by [59], and the case of $m=0$ was considered by [52]. Secondly, proper distributions of the squared ReLU function remain unknown. Incorporating it at any layer of a DNN with a depth of $L$ might lead to a polynomial degree of $2^{L}$ with respect to parameters, which poses challenges in training. Thirdly, the optimality of the approximation process remains unclear due to the complexity associated with higher-order derivatives of DNNs using ReLU and squared ReLU activations. To establish optimality, estimating the optimal Vapnik-Chervonenkis dimension (VC-dimension) and pseudo-dimension [2, 56, 1, 47] of their higher order derivatives, which characterize the complexity or richness of a function set, need to be estimated. Even for DNNs without derivatives, as demonstrated in [3], obtaining optimal bounds is challenging. The upper VC-dimension and pseudo-dimension for DNNs using squared ReLU is $\mathcal{O}(N^{2}L^{3}\log_{2}N\log_{2}L)$ or $\mathcal{O}(N^{3}L^{2}\log_{2}N\log_{2}L)$. However, this result is not directly applicable to proving the optimality of squared ReLU-based DNN approximation due to a polynomial gap in depth $L$.
@@
In this paper, the deep super ReLU networks (DSRNs) is presented to address these three questions, as illustrated in Fig. 1. The DSRN architecture consists of two components for a network with depth $L$: ReLU DNNs with a depth of approximately $\mathcal{O}(L)$, and the square of ReLU DNNs with a depth of approximately $\mathcal{O}(\log_{2}L)$—significantly shallower than the first component.
@@
In terms of training, our DSRNs predominantly employ ReLU activations, with the square of ReLU appearing only at the end of the DNNs. This enhances the smoothness of DNNs while preserving a straightforward network structure without intricate composition formulas. This contrasts with [59], which did not specify where to apply the square of ReLU, leading to training difficulties. Our experiments in Sec. 7 demonstrate that our proposed DNN (DSRN) trains effectively, unlike those using the square of ReLU throughout, which are difficult to train due to gradient vanishing. For the training of DSRN, we discuss more in Section 1.1 later.
Finally, we estimate the generalization error of DSRNs in Theorem 14 under Sobolev training [8]. Although there are works focusing on the generalization error for Sobolev training—such as for physics-informed neural networks [37, 29, 42, 10] and the Deep Ritz method [30, 38, 59, 31]—most existing results consider continuous approximators. In these cases, the approximation rate is consistent with classical results [12, 13], where the parameter size is uniformly bounded. Under this setting, the generalization error can be bounded by directly estimating the covering number of the parameter space.
However, in the generalization analysis of deep neural networks, if we aim to achieve super-convergence—as shown in Theorem 1—many parameters can become large and are not uniformly bounded. Therefore, a different approach is required to estimate the generalization error in Sobolev training when the network achieves super-convergence. Our results are nearly optimal with respect to the number of sample points and significantly improve upon related works such as [59, 30]. In Theorem 14, the obtained generalization rate is nearly optimal with respect to the number of sampling points. There are two key reasons for this improvement. First, we employ covering number arguments instead of Rademacher complexity to bound the generalization error. Rademacher complexity is a global measure that may lose local structural information, which can result in suboptimal error rates. To achieve optimal generalization error bounds, it is necessary to adopt localized complexity measures, such as local Rademacher complexity [4, 60] or covering number techniques [23, 35]. Both approaches can yield optimal bounds, in this work, we adopt the covering number approach. The second reason lies in our tight estimates of the VC-dimension and pseudo-dimension of higher-order derivatives of deep neural networks. These estimates lead to sharper covering number bounds. To the best of our knowledge, this is the first work that explicitly analyzes the VC- and pseudo-dimensions of higher-order derivatives in neural networks, and applies them directly to the error analysis in Sobolev training.
1.1 Training of DSRN
In this section, we discuss the training of DSRNs, which naturally leads to the following fundamental questions. A DSRN, defined as the composition of a squared ReLU neural network with another ReLU neural network, is generally non-smooth for most parameter choices. This raises two important questions:
Questions: (i) Can such a neural network approximate a target function measured in a Sobolev norm? (ii) If so, can this type of neural network be effectively trained in practice?
Answer to Question (i) and proof sketch.
Theorem 2 together with Corollary 8 gives an affirmative answer. We briefly outline the construction.
(1) Local polynomial approximation. By the Bramble–Hilbert lemma [6], on each cell of a shape-regular partition of the domain there exists a polynomial (of the prescribed order) that approximates the target function in the relevant Sobolev norm with the standard rate.
(2) Exact realization of polynomials by squared-ReLU. Squared-ReLU networks can represent polynomials exactly; more specifically, a degree-$n$ polynomial can be realized by a squared-ReLU network of depth $\mathcal{O}(\log_{2}n)$ [25]. Hence, on each cell we implement the local polynomial by a very shallow squared-ReLU subnetwork (serving as a smooth “feature extractor”).
(3) Gating and global assembly. To “turn on” the appropriate local feature, we use a (standard) ReLU network to implement piecewise-constant gating weights; an efficient construction follows the bit-extraction paradigm [3]. This yields a global piecewise model. The only nonsmoothness arises along the interfaces where neighboring local pieces interact (the “bad regions”).
(4) Smoothing via partitions of unity within the architecture. We build several such subnetworks using differently shifted local partitions so that their interface sets do not coincide. We then combine them with a smooth partition of unity (PoU) whose supports are chosen to avoid the bad region of at least one subnetwork on each point of the domain. Because DSRNs can implement $(x)_{+}^{s}$ for any $s\in\mathbb{N}$, they can represent truncated-power bases and hence B-splines; consequently, smooth PoUs (and smooth piecewise-polynomial functions) can be realized within the DSRN [9]. This yields the desired Sobolev-norm approximation claimed in Theorem 2 and Corollary 8.
Further mathematical details can be found in Section 4.1. The overall structure of the neural network is illustrated in Fig. 2.
Figure 2: The architecture of the smooth DSRN.
For the second question, we divide the discussion into two parts.
First, can we actually run the code? The answer is yes. Although the initial guess of a neural network may be non-smooth, the non-smooth points lie on a set of measure zero in the domain, which is unlikely to be sampled. Even if some points are sampled at these locations, we can simply resample during training to avoid them. Therefore, the experiments are indeed feasible.
Second, can such neural networks learn functions and match higher-order derivatives? The answer is also affirmative. During training, the DSRN gradually becomes smooth. After random initialization, the DSRN outputs exhibits jump discontinuities, leading to highly irregular derivatives. However, since the loss function involves high-order derivatives, the training process gradually reduces these discontinuities. The theoretical analysis guarantees that there exist globally smooth DSRNs capable of approximating target functions in Sobolev norms. Consequently, the network systematically eliminates jump discontinuities and converges toward smooth function approximations with respect to higher-order Sobolev norms.
In practice, there may still remain very small jump points, but their effect is negligible, and the network can already approximate the derivatives of the target function well. If even higher smoothness is desired, one can resample around these jump points to further reduce the gaps.
To illustrate this dynamic smoothing process, we use a DSRN to approximate $f(x)=x^{3}$ and Fig. 3 shows the evolution of the second-order derivative estimated by DSRN during the training. The learned second-order derivatives quickly approach the ground truth $f^{\prime\prime}(x)=6x$ across the domain, demonstrating that DSRN progressively captures the smoothness of the target function in terms of the Sobolev norm. More detailed experiments for solving PDEs can be found in Sec. 7.
Figure 3: Evolution of the second-order derivative estimated by DSRN during training. The panels (arranged from top to bottom and left to right) correspond to the initial state, iterations 10,000, 20,000, 30,000, 40,000, and the final iteration.
1.2 Main contributions
@@
$\bullet$ We provide an upper bound estimation for the pseudo-dimension of high-order derivatives of DNNs. By utilizing this estimation, we can obtain the nearly optimal generalization error of DNNs in Sobolev spaces that contain high-order derivatives of DNNs.
@@ 2.1 Neural networks
1. Matrices are denoted by bold uppercase letters. For example, $\bm{A}\in\mathbb{R}^{m\times n}$ is a real matrix of size $m\times n$ and $\bm{A}^{\intercal}$ denotes the transpose of $\bm{A}$.
2. Vectors are denoted by bold lowercase letters. For example, $\bm{v}\in\mathbb{R}^{n}$ is a column vector of size $n$. Furthermore, denote $\bm{v}(i)$ as the $i$-th elements of $\bm{v}$.
3. For a $d$-dimensional multi-index $\bm{\alpha}=[\alpha_{1},\alpha_{2},\cdots\alpha_{d}]\in\mathbb{N}^{d}$, we denote several related notations as follows:
$(a)~|\bm{\alpha}|=\left|\alpha_{1}\right|+\left|\alpha_{2}\right|+\cdots+\left|\alpha_{d}\right|;$
$(b)~\bm{x}^{\bm{\alpha}}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{d}^{\alpha_{d}},~\bm{x}=\left[x_{1},x_{2},\cdots,x_{d}\right]^{\intercal};$
$(c)~\bm{\alpha}!=\alpha_{1}!\alpha_{2}!\cdots\alpha_{d}!.$
4. Assume $\bm{n}\in\mathbb{N}_{+}^{m}$, and $f$ and $g$ are functions defined on $\mathbb{N}_{+}^{m}$, then $f(\bm{n})=\mathcal{O}(g(\bm{n}))$ means that there exists positive $C$ independent of $\bm{n},f,g$ such that $f(\bm{n})\leq Cg(\bm{n})$ when all entries of $\bm{n}$ go to $\infty$.
5. Define $\sigma_{1}(x):=\max\{0,x\}$ and $\sigma_{2}:=\sigma_{1}^{2}(x)$. We call the neural networks with activation function $\sigma_{t}$ with $t=i$ as $\sigma_{i}$ neural networks ($\sigma_{i}$-NNs), $i=1,2$. With the abuse of notations, we define $\sigma_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ as $\sigma_{i}(\bm{x})=\left[\begin{array}[]{c}\sigma_{i}(x_{1})\\ \vdots\\ \sigma_{i}(x_{d})\end{array}\right]$ for any $\bm{x}=\left[x_{1},\cdots,x_{d}\right]^{T}\in\mathbb{R}^{d}$.
@@
$\bm{x}=\tilde{\bm{h}}_{0}\stackrel{{\scriptstyle W_{1},b_{1}}}{{\longrightarrow}}\bm{h}_{1}\stackrel{{\scriptstyle\sigma_{i}}}{{\longrightarrow}}\tilde{\bm{h}}_{1}\ldots\stackrel{{\scriptstyle W_{L},b_{L}}}{{\longrightarrow}}\bm{h}_{L}\stackrel{{\scriptstyle\sigma_{i}}}{{\longrightarrow}}\tilde{\bm{h}}_{L}\stackrel{{\scriptstyle W_{L+1},b_{L+1}}}{{\longrightarrow}}\phi(\bm{x})=\bm{h}_{L+1},$
where $\bm{W}_{i}\in\mathbb{R}^{N_{i}\times N_{i-1}}$ and $\bm{b}_{i}\in\mathbb{R}^{N_{i}}$ are the weight matrix and the bias vector in the $i$-th linear transform in $\phi$, respectively, i.e., $\bm{h}_{i}:=\bm{W}_{i}\tilde{\bm{h}}_{i-1}+\bm{b}_{i},~\text{ for }i=1,\ldots,L+1$ and $\tilde{\bm{h}}_{i}=\sigma_{i}\left(\bm{h}_{i}\right),\text{ for }i=1,\ldots,L.$ In this paper, an DNN with the width $N$ and depth $L$, means (a) The maximum width of this DNN for all hidden layers is less than or equal to $N$. (b) The number of hidden layers of this DNN is less than or equal to $L$.